OpenStudy (anonymous):

Differentiate using first principle , f(x)=2/x

OpenStudy (kc_kennylau):

\[\Large f'(x)=\lim_{h\rightarrow0}\frac{f(x+h)-f(x)}{h}=\lim_{h\rightarrow0}\frac{\frac2{x+h}-\frac2x}h\]

OpenStudy (kc_kennylau):

Can you continue from this? :)

OpenStudy (isaiah.feynman):

 Being so long since I looked at this this!